OpenStudy (anonymous):

The GPE of a 2.0 kg object that is 4.0 m high is ...?

OpenStudy (anonymous):

8.0 J. 78.4 J. 6.0 J. 49.8 J. ?

OpenStudy (ash2326):

GPE=M*g*h M=mass g=acceleration due to gravitational force 9.8m/s^2 h=height GPE=2*9.8*4=78.4 J option b